# Carlo Orientale Caputo SISSA - International School for Advanced Studies, 34136 Trieste, Italy Elias Seiffert

### v3

@@ Abstract
Abstraction is the process of extracting the essential features from raw data while ignoring irrelevant details. This is similar to the process of focusing on large-scale properties, systematically removing irrelevant small-scale details, implemented in the renormalisation group of statistical physics. This analogy is suggestive because the fixed points of the renormalisation group offer an ideal candidate of a truly abstract – i.e. data independent – representation.
@@
It has been observed that abstraction emerges with depth in neural networks. Deep layers of neural network capture abstract characteristics of data, such as ”cat-ness” or ”dog-ness” in images, by combining the lower level features encoded in shallow layers (e.g. edges). Yet we argue that depth alone is not enough to develop truly abstract representations. We advocate that the level of abstraction crucially depends on how broad the training set is. We address the issue within a renormalisation group approach where a representation is expanded to encompass a broader set of data. We take the unique fixed point of this transformation — the Hierarchical Feature Model – as a candidate for an abstract representation. This theoretical picture is tested in numerical experiments based on Deep Belief Networks trained on data of different breadth. These show that representations in deep layers of neural networks approach the Hierarchical Feature Model as the data gets broader, in agreement with theoretical predictions.
Following Marr [1], one may argue that the conceptual underpinnings of cognitive functions are independent of whether they are implemented in-silico or in a biological brain. In this spirit, this paper adopts artificial neural networks (ANN) as a playground to explore how abstraction arises from the process of making sense of raw complex data. To illustrate this, consider vision. Both in biological brains [1] and in ANNs [2], vision involves a hierarchical organization of representations: shallow layers detect low-level features, such as edges [3], while deeper layers integrate these features to recognize more abstract, higher-order constructs like objects and faces [2]. In particular, deeper layers are capable of recognising an object or a face irrespective of it’s position, orientation, scale, or of context [^1] This ability can be promoted in ANN by either augmenting the data using invariances [4] or by explicitly implementing them in their architecture – as in convolutional neural networks [2]. Yet even simple neural networks are able to develop a convolutional structure by themselves [5]. [6, 7]. This parallel underscores the broader principle that abstraction emerges through layered processing, regardless of the underlying substrate on which the process is implemented.
In general, abstract representations extracted from data have been discussed in terms of “cognitive maps” [8, 9]. A cognitive map is not only an efficient and flexible scaffold of data, but it is also endowed with a structure of relations – uncovered from the data – that enables abstract computation [^2] Relational structures such as “Alice is the daughter of Jim” and ”Bob is Alice’s brother” allow for computations (e.g. ”Jim is Bob’s father”) which are invariant with respect to the context (Alice, Jim and Bob can be replaced by any triplet of persons that stand in the same relation, in this example) [8]. [8, 9] and supports complex functions. For example, spatial navigation relies on the representation built by several assemblies of specialised neurons, such as grid cells [10, 11].
At even higher levels of cognition, representations should integrate data from a broader set of domains, or perceptual modalities [12], each of which may be organised according to cognitive maps of a different nature. For example, while visual stimuli are described by object manifolds with supposedly local euclidean topology [13, 14], odours have been suggested to be organised in hyperbolic spaces [15]. Higher order representations that integrate the two should therefore be even more abstract, i.e. independent of the data.
A lot has been understood about the role of depth in learning [^3] Strictly speaking, most of these insights pertain to supervised learning, yet we assume they reveal properties that are relevant also for unsupervised learning, which is our focus. [13, 16, 17, 14, 18, 19, 20]. In particular, depth exploits the compositional structure of the data boosting training performance [16, 18, 20] and classification capacity [17]. Indeed, inner layers of ANN portray data that correspond to the same object as “object manifolds” [13] that become better and better separable with depth [17], while extracting hierarchies of features that promote taxonomic abstraction [^4] Taxonomic abstraction is based on the idea that objects are similar if they share the same features (e.g. ”cats” and ”dogs” both have four legs, a tail, etc) and belong to the same category (mammals). Taxonomic abstraction is fundamentally distinct from thematic abstraction, which is based on co-occurrence between objects that share no features (e.g. ”cat” and ”sofa”), as discussed in [21]. In deep neural networks [18, 20] shallow layers capture statistical associations (thematic) while deep layers encode compositional, taxonomic structures. A similar transition is observed in humans with development: while children favour thematic abstractions, adults more frequently rely on taxonomic (or categorical) structures [22]..
A possible road to abstraction has been suggested [23, 24] exploiting the analogy between the processing of data in deeper and deeper layers and the renormalisation group (RG) in statistical physics [25] (see Fig. 1 A). This analogy is based on the observation that higher order features in learning are akin to large scale properties in statistical models, which are those that the RG singles out. This idea is suggestive for at least two reasons: First the RG is the theoretical tool to study critical phenomena [25] and both artificial and biological learning exhibit critical features [26, 27, 28, 29, 30, 31]. Second, and most importantly, repeated RG transformations lead to universal distributions, which are an ideal candidate for abstract representations.
In this paper we argue that this analogy misses a fundamental ingredient, which is breadth, i.e. the diversity in the input data or stimuli [^5] We distinguish breadth from width, a term commonly used in the literature to denote the number of variables in different layers.. It is common sense that a representation encompassing a broader universe of circumstances should be more abstract than one describing only a limited domain. Likewise, we argue that representations in higher levels of the cognitive hierarchy, that integrate a broader set of stimuli, should be more abstract, i.e. independent of the data. This paper approaches the problem of characterising such abstract representations on the basis of their sole statistical properties, with no reference to what is being represented. In this respect, we depart from the typical “tuning curve” approach in computational neuroscience, in which levels of abstraction are assessed in terms of the features of the data – e.g. edges or faces – that a representation encodes, as well as from other efforts aimed at finding those properties that make structured data learnable (see e.g. [18, 20, 32]). Focusing on the marginal distribution of activations of inner layers allows us to characterise abstract representations purely through information-theoretic principles. We focus on the simple case of static data: a representation in this paper is a probability distribution over a set of binary variables.
@@
In this setting, we show that abstract representations emerge as the fixed point of a RG transformation whereby a representation for a given domain is updated to describe data from a broader domain. In this process, assuming a constant coding cost, low level details are sacrificed in order to make space for high level features describing the organisation of the data within the broader domain. Fig. 1 C sketches this process for an illustrative example. This process of zooming out to a broader domain while loosing low level details can also be inverted, by zooming into a specific part of the data, thus uncovering low level details (see Fig. 1 B). We show that in both cases, the transformation has a unique fixed point that coincides with the Hierarchical Feature Model (HFM), recently introduced in [33]. This is reassuring for at least two reasons: First the HFM is a maximum entropy model fully determined by a single sufficient statistics, which is the average level of detail of the features, or the coding cost. This is indeed the only relevant variable in an abstract representation. Second, the HFM satisfies the principle of maximal relevance. The relevance has been recently introduced [34] as a quantitative measure of “meaning” that captures Barlow’s intuition [35] that meaning is carried by redundancy. We refer to Section 1.1 for a brief discussion of the relevance, or to Ref. [36] for an extended account. Let it suffice to say that internal representations of well trained learning machines have been shown to satisfy the principle of maximal relevance [37, 38].
While a transformation in terms of breadth has a unique fixed point, we shall also argue that a similar transformation in terms of depth alone may have many fixed points. This is an important observation for the second part of the paper, where we shall search for evidence of our theory in empirical data of deep belief networks trained on different datasets. Our numerical experiments, presented in Section 3, will show that convergence to the abstract representation of the HFM requires the combined effect of depth and breadth. Section 4 discusses the results and provides some concluding remarks. All technical details are relegated to the Appendix [^6] The present paper supersedes the preliminary results presented in the Master thesis of one of us [39]..
@@ 1 The framework
In this paper, a representation is a probability distribution $p(\mathbf{s})$ over a string of binary variables $\mathbf{s}=(s_{1},\ldots,s_{n})$ ($s_{i}\in\{0,1\}$), that model the activation of “neurons” inside a learning machine. For example, $\mathbf{s}=\mathbf{s}^{(\ell)}$ can be the variables in the $\ell^{\rm th}$ layer of a neural network with many layers. We shall focus on unsupervised learning and generative models such as a Deep Belief Network (DBN), that mathematically can be seen as a joint probability distribution
$p(\mathbf{x},\mathbf{s}^{(1)},\ldots,\mathbf{s}^{(L)})=p(\mathbf{x}|\mathbf{s}^{(1)})p(\mathbf{s}^{(1)}|\mathbf{s}^{(2)})\cdots p(\mathbf{s}^{(L-1)}|\mathbf{s}^{(L)})p(\mathbf{s}^{(L)})$ (1)
of the variables $\mathbf{x}=(x_{1},\ldots x_{m})$ in the visible layer and the variables $\mathbf{s}^{(\ell)}=(s_{1},\ldots,s_{n_{\ell}})$ in the $L$ hidden layers ($\ell=1,\ldots,L$). Eq. (1) also highlights the Markovian nature of information processing whereby the activation of layer $\ell$ only depends on data through layer $\ell-1$. We focus on this probability distribution after the network has been trained on some data $\hat{\mathbf{x}}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})$ by maximising the log-likelihood $\mathcal{L}(\hat{\mathbf{x}})=\sum_{i}\log p(\mathbf{x}_{i})$ over the parameters of the network. We refer to Appendix A for a more detailed discussion of such networks. For our purposes, let it suffice to say that the object of our study will be the marginal distribution $p(\mathbf{s}^{(\ell)})$ of $\mathbf{s}^{(\ell)}$ after training. In a well trained network, sampling $\mathbf{s}^{(\ell)}$ from this distribution and propagating the state until the visible layer generates data points which are statistically indistinguishable from the data.

### v4

# Absolute abstraction: a renormalisation group approach

@@ Abstract
Abstraction is the process of extracting the essential features from raw data while ignoring irrelevant details. It is well known that abstraction emerges with depth in neural networks, where deep layers capture abstract characteristics of data by combining lower level features encoded in shallow layers (e.g. edges). Yet we argue that depth alone is not enough to develop truly abstract representations. We advocate that the level of abstraction crucially depends on how broad the training set is. We address the issue within a renormalisation group approach where a representation is expanded to encompass a broader set of data. We take the unique fixed point of this transformation — the Hierarchical Feature Model – as a candidate for a representation which is absolutely abstract. This theoretical picture is tested in numerical experiments based on Deep Belief Networks and auto-encoders trained on data of different breadth. These show that representations in neural networks approach the Hierarchical Feature Model as the data get broader and as depth increases, in agreement with theoretical predictions.
Abstraction refers to the ability to represent general concepts from sensory input or from raw data. It is well known that representations which are increasingly independent of details arise when raw data is processed in a hierarchy of deeper and deeper layers – both in-silico [1] and in-vivo [2]. Here we argue that universal representations emerge when the process of removing irrelevant details is iterated indefinitely on a universe of data that simultaneously expands in variety, that is when increasing depth is combined with expanding breadth [^1] We distinguish breadth from width, a term commonly used in the literature to denote the number of variables in different layers.. These universal representations encode the concept of absolute abstraction.
We address this in a minimal setting of unsupervised learning from static data (e.g. images). We shall focus on the internal representation [^2] A representation in this paper is a probability distribution over a set of binary variables. of the data in a deep layer of a learning machine and on how it adapts when the training data expands in scope. This process can be cast within a renormalisation group (RG) framework that allows us to identify the fixed point of the RG transformation with the ultimate outcome of this process when it is repeated indefinitely. The analogy between the process of learning broader and broader data in deeper and deeper leyers and the RG in statistical physics [3] (see Fig. 1 A) is based on the observation that higher order features in learning are akin to large scale properties in statistical models, which are those that the RG singles out. This idea is corroborated by similarities between coarse graining and the evolution of representations with depth [4, 5].
The fact that representations converge to a universal, data-independent, distribution is consistent with the expectation that a representation encompassing a broader universe of circumstances should be more abstract – i.e. data independent – than one describing only a limited domain. In this picture, all data-specific information is transferred to the parameters through which the activity of deep layers propagates to the visible layer. In other words, the fixed point of the RG describes how data is organised in the ideal limit of infinite depth and breadth, irrespective of what the data is about.
Hence this paper approaches the problem of characterising abstract representations on the basis of their sole statistical properties, with no reference to what is being represented, guided only by information-theoretic principles.
@@
When a representation with a fixed capacity is updated to describe data from a broader domain, low level details need to be sacrificed in order to make space for high level features describing the organisation of the data within the broader domain. Fig. 1 C sketches this process for an illustrative example. This process of zooming out to a broader domain while loosing low level details can also be inverted, by zooming into a specific part of the data, thus uncovering low level details (see Fig. 1 B). We show that in both cases, the corresponding RG transformation has a unique fixed point which is related to the Hierarchical Feature Model (HFM), recently introduced in [6]. This is reassuring for at least two reasons: First the HFM is a maximum entropy model fully determined by a single sufficient statistics, which is the average level of detail of the features, or the coding cost. This is indeed the only relevant variable in an abstract representation. Second, the HFM satisfies the principle of maximal relevance [^3] The relevance has been recently introduced [7] as a quantitative measure of “meaning” that captures Barlow’s intuition [8] that meaning is carried by redundancy. We refer to Appendix A.1 for a brief discussion of the relevance, or to Ref. [9] for an extended account., a principle that has been suggested to characterize most informative representations and that also well-trained learning machines have been shown empirically to obey [10, 11].
The rest of the paper is organised as follows: The next Section places our contribution within the broader literature on machine learning and neuroscience. Then, in Section 2 we lay out the framework for our analysis. The RG analysis is presented in Section 3. Section 4 discusses numerical experiments on Deep Belief Networks and auto-encoders that corroborate the theoretical predictions. Within their limited expressive capacity, the networks we studied show that representations in deep layers approach the HFM under the combined effects of depth and breadth. Section 5 discusses the results and provides some concluding remarks. All technical details are relegated to the Appendix [^4] The present paper supersedes the preliminary results presented in the Master thesis of one of us [12]..
1 Literature review
Following Marr [2], we argue that the conceptual underpinnings of the capacity of abstraction are independent of its algorithmic implementation or of whether it is implemented in-silico or in a biological brain. This means that both cognitive neuroscience and machine learning may provide useful insights to characterize abstraction.
Vision provides the paradigmatic case for exploring how abstract representations arise. Both in biological brains [2] and in artificial neural networks (ANNs) [1], vision involves a hierarchical organization of representations: shallow layers detect low-level features, such as edges [13], while deeper layers integrate these features to recognize more abstract, higher-order constructs like objects and faces [1]. In particular, deeper layers are capable of recognising an object or a face irrespective of it’s position, orientation, scale, or of context [^5] This ability can be promoted in ANN by either augmenting the data using invariances [14] or by explicitly implementing them in their architecture – as in convolutional neural networks [1]. Yet even simple neural networks are able to develop a convolutional structure by themselves [15]. [16, 17]. This parallel underscores the broader principle that abstraction emerges through layered processing, regardless of the underlying substrate on which the process is implemented. Yet it has also been argued [18] that the limit of infinite depth is singular, because too much depth without constraints washes away meaningful structure in the input data.
In general, abstract representations extracted from data have been discussed in terms of “cognitive maps” [19, 20]. A cognitive map is not only an efficient and flexible scaffold of data, but it is also endowed with a structure of relations – uncovered from the data – that enables abstract computation [^6] Relational structures such as “Alice is the daughter of Jim” and ”Bob is Alice’s brother” allow for computations (e.g. ”Jim is Bob’s father”) which are invariant with respect to the context (Alice, Jim and Bob can be replaced by any triplet of persons that stand in the same relation, in this example) [19]. [19, 20] and supports complex functions. For example, spatial navigation in rats relies on the representation built by several assemblies of specialised neurons, such as grid cells [21, 22].
At the highest levels of cognition, representations should integrate data from a broad set of domains, or perceptual modalities [23], each of which may be organised according to cognitive maps of a different nature. For example, while visual stimuli are described by object manifolds with supposedly euclidean topology [24, 25], odours have been suggested to be organised in hyperbolic spaces [26]. Higher order representations that integrate the two should therefore be even more abstract, i.e. independent of the data.
A lot has been understood about the role of depth in learning [^7] Strictly speaking, most of these insights pertain to supervised learning, yet we assume they reveal properties that are relevant also for unsupervised learning, which is our focus. [24, 27, 28, 25, 29, 30, 31]. In particular, depth exploits the compositional structure of the data boosting training performance [27, 29, 31] and classification capacity [28]. Indeed, inner layers of ANN portray data that correspond to the same object as “object manifolds” [24] that become better and better separable with depth [28], while extracting hierarchies of features that promote taxonomic abstraction [^8] Taxonomic abstraction is based on the idea that objects are similar if they share the same features (e.g. ”cats” and ”dogs” both have four legs, a tail, etc) and belong to the same category (mammals). Taxonomic abstraction is fundamentally distinct from thematic abstraction, which is based on co-occurrence between objects that share no features (e.g. ”cat” and ”sofa”), as discussed in [32]. In deep neural networks [29, 31] shallow layers capture statistical associations (thematic) while deep layers encode compositional, taxonomic structures. A similar transition is observed in humans with development: while children favour thematic abstractions, adults more frequently rely on taxonomic (or categorical) structures [33].. It has been proposed that the superior cognitive abilities of homo-sapiens crucially rely on the expansion of the neo-cortex in depth and to specific mutations that supported reliable neural computation [34]. In this respect, our results suggest that exposure to a broad variety of data and stimuli is also essential for the emergence of abstract representations that may support intelligence.
The analogy between the processing of data in deeper and deeper layers and the renormalisation group (RG) in statistical physics [3] has been suggested [4, 5, 35] long ago. This analogy is suggestive because the RG is the theoretical tool to study critical phenomena [3] and both artificial and biological learning exhibit critical features [36, 37, 38, 39, 40, 41]. Koch et al. [5] have shown that successive training in deeper layers of neural networks performs an operation similar to coarse graining in the RG. Yet the RG transformation does not only involve coarse graining. It also involves rescaling, which is the operation by which the size of the coarse grained system is restored (see Fig. 1 A). We argue that this ingredient corresponds to a change in breadth, i.e. in the diversity in the input data or stimuli. It is indeed common sense that a representation encompassing a broader universe of circumstances should be more abstract than one describing only a limited domain. Likewise, we argue that representations in higher levels of the cognitive hierarchy, that integrate a broader set of stimuli, should be more abstract, i.e. independent of the data.
The emphasis on a data-independent characterisation of abstraction based only of statistical properties contrasts with the “tuning curve” approach [13, 42], in which levels of abstraction are assessed in terms of the features of the data – e.g. edges or faces – that a representation encodes. Likewise, although we implicitly assume that the representations we are interested in describe structured and learnable data, we shall refrain from assuming any specific structure like the ones proposed in Refs. [29, 43, 31].
@@ 2 The framework
In this paper, a representation is a probability distribution $p(\mathbf{s})$ over the configurations $\mathbf{s}$ of a hidden layer of a generative neural network, such as a Deep Belief Network (DBN) [44] or an auto-encoders [45]. Hence we focus on unsupervised learning and we assume that the network has been trained on some data $\hat{\mathbf{x}}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})$ with a non-trivial structure. The representation $p(\mathbf{s})$ is the marginal distribution of activation of variables in the hidden layer induced by the distribution $p(\mathbf{x})$ of the data, i.e.
$p(\mathbf{s})=\sum_{\mathbf{x}}p(\mathbf{s}|\mathbf{x})p(\mathbf{x})\,.$ (1)
We shall confine our discussion to the case where $\mathbf{s}_{1:n}=(s_{1},\ldots,s_{n})$ is a string of binary variables $s_{i}=0,1$. Otherwise, we shall keep our discussion as general as possible for the time being. More details on network architectures and the data will be given in Section 4 and Appendix C.2, where we shall test our predictions on specific models.